TRANSITION DIPOLE MOMENT AND LINE STRENGTHS FOR THE ν_{2} FUNDAMENTAL BAND OF THE METHYL RADICAL

G. D. STANCU AND J. RÖPCKE, INP-Greifswald, 17489 Greifswald, Germany; \underline{P. B. DAVIES}, Department of Chemistry, University of Cambridge, Cambridge, U.K..

The determination of methyl radical concentrations in terrestrial and astronomical sources using infrared spectroscopy relies on the availability of accurate line strengths and transition dipole moments. The ν_{2} fundamental of CH_{3} near 606 cm^{−1} is particularly useful for this purpose but the current value of its transition dipole moment is uncertain by at least 20 %. We have measured the line strength of 9 Q-branch lines of the 2_{0}^{1} band of CH_{3} from Q(1,1) to Q(12,12) using diode laser absorption spectroscopy. The method is based on following the recombination rate of methyl radicals in a pulsed discharge in tertiary butyl peroxide heavily diluted in argon^{a}. The translational, rotational and vibrational temperatures of the radical were precisely determined from line widths and relative line intensities in both the ν_{2} fundamental and hot bands. The new value of the transition dipole moment is 0.22±0.02 D which is compatible with most previous measurements and calculations, and with the transition dipole moment of the ν_{3} band.

aC.Yamada and E. Hirota, J.Chem. Phys. \underline{78}, (2) 669 (1983).